# On automorphism groups of smooth hypersurfaces

### v1

@@ 7. Proof of the main theorem
In this section, using Theorem 6.1 and the refined bounds in Section 5, we classify smooth forms with large imprimitive or reducible automorphism groups (Theorems 7.1 and 7.4). Combining these results, we prove a stronger form (Theorem 7.5) of Theorem 1.1.
@@ 7.1. Imprimitive ${\rm Aut}(F)$
By the assumption, $l({\rm Aut}(F))$ is of exponential type $r^{k}$ with $r\geq 2$ and $k\geq 2$, and ${\rm Aut}(F)$ has a primitive-constituent sequence $\mathcal{H}:=(H_{1},\ldots,H_{k})$ with $H_{i}\subset{\rm PGL}(r,{\mathbb{C}})$ and $H_{i}\cong H_{j}$ for $1\leq i,j\leq k$. By Lemma 4.6, we have $R(\mathcal{H},d)\geq 1$. By direct calculation (see (4.1)), we get $|H_{i}|>d^{r-1}\cdot r!$. It suffices to consider two cases: $r\geq 3$ and $r=2$.
Case $r\geq 3$. If $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$, then by Theorem 6.1, we have $(r,d,|H_{1}|)=(3,4,168)$, $(3,6,360)$, $(4,4,1920)$, $(6,6,6531840)$. By direct calculation, we have $R(\mathcal{H},d)<1$, a contradiction. Thus, $H_{1}$ preserves no smooth forms of degree $d$. Since ${\rm JC}(r)>106$ for $r\geq 3$, by Lemmas 5.1, 5.5 and 4.14, we get $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$, a contradiction.
Case $r=2$. If $H_{1}$ preserves no smooth forms of degree $d$, then $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$ by similar arguments as in case $r\geq 3$. Thus, we may assume $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$. By Lemma 6.4 , if $H_{1}\cong A_{4},S_{4},A_{5}$, then $H_{1}$ preserves no smooth forms in $2$ variables of degree less than $4,6,12$ respectively. If $H_{1}\cong A_{4}$, then $d\geq 4$ and $R(\mathcal{H},d)<1$, a contradiction. If $H_{1}\cong S_{4}$, then $d\geq 6$ and $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,6)$. Then $R(\mathcal{H},d)=4/3$ and by Lemma 5.6, $F$ can be expressed as $F_{1}(x_{1},x_{2})+F_{2}(x_{3},x_{4})$, where $F_{i}$ are preserved by $H_{i}$. Therefore, $F$ is unique up to isomorphism by Lemma 6.4. If $H_{1}\cong A_{5}$, then $d=12$ or $d\geq 20$ by Lemma 6.4. So $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,12),(3,12)$. Similar to the case $H_{1}=S_{4}$, we have the uniqueness of $F$. ∎
@@ 7.2. Reducible ${\rm Aut}(F)$
By direct calculations, if $n_{m}=1$ or $m\geq 3$ or $n_{1}>2$, then all fractions on the right side are less than $1$, a contradiction to $q\geq 1$. Thus, we have $m=2$, $n_{1}=n_{2}=2$ and $q=\frac{25}{24}$. ∎
@@
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 3$. If ${\rm Aut}(F)$ is reducible, then $|{\rm Aut}(F)|<d^{r}\cdot r!$
@@
To prove the theorem by contradiction, we suppose $\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\geq 1$. Let $G:={\rm Aut}(F)$. By Theorem 3.9, we may assume at least one subdegree of $G$ is larger than $1$. Let $V_{i},W_{ij},k_{i},r_{ij},H_{ij},m,F_{ij}$ be as in $\bf{Set\textrm{-}up}$ 5.3. We define $\beta$ be to $1$ (resp. $2$) if $l(G)$ has at least two different subdegrees (resp. all subdegrees of $G$ are the same). From the definitions, we have
@@
for any pair $(i,j)$ and any $1\leq k\leq k_{i}$. By Lemmas 4.6 and 4.14, we obtain $R(l(G),d)<60\beta$. If there is a monomial
@@
Case (1). Note that $V_{i}$ $(i=1,\ldots,m_{1})$ are irreducible $G$-stable subspaces of dimension $r_{i}:={\rm dim}(V_{i})=r_{i1}k_{i}$ and the irreducible linear groups $G_{i}:=p_{i}(G)$ preserve the smooth forms $F_{i}$ of degree $d$. Here $p_{i}:G\rightarrow{\rm GL}(V_{i})$ is given by $g\mapsto g|V_{i}$. Let $V^{\prime}:=\oplus_{i>m_{1}}V_{i}$ and $p^{\prime}:G\rightarrow{\rm GL}(V^{\prime})$ be given by $g\mapsto g|V^{\prime}$. Let $G_{m_{1}+1}:=p^{\prime}(G)$. Then $r=r_{1}+\cdots r_{m_{1}}+r_{m_{1}+1}$, where $r_{m_{1}+1}:={\rm dim}(V^{\prime})$. By Proposition 7.2, we have $\mathfrak{R}_{F_{i}}=\frac{|G_{F_{i}}|}{d^{r_{i}}\cdot r_{i}!}\leq 5/2$ $(i=1,2,\ldots,m_{1})$. By Theorem 3.9, we have $\mathfrak{R}_{F_{m_{1}+1}}=\frac{|G_{F_{m_{1}+1}}|}{d^{r_{m_{1}+1}}\cdot r_{m_{1}+1}!}\leq 1$. Clearly $G_{i}\subset G_{F_{i}}$ $(i=1,2,\ldots,m_{1}+1)$. Then by Lemma 7.3 and
$1\leq\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\leq\frac{|G_{1}|\cdot|G_{2}|\cdots|G_{m_{1}}|\cdot|G_{m_{1}+1}|}{d^{r}\cdot r!}\leq\frac{\prod_{i=1}^{m_{1}+1}\mathfrak{R}_{F_{i}}\cdot r_{i}!}{(r_{1}+\cdots+r_{m_{1}+1})!},$
@@
If $F^{\prime}$ is zero, then ${\rm Aut}(F)$ is irreducible (see Theorem 7.1), a contradiction. If $F^{\prime}$ is not zero, then by Lemmas 5.6 and 6.4, we get $\mathfrak{R}_{F}<1$, a contradiction. This proves case (1).
Case (2). Let $\mathcal{A}$ and $c$ be as in Lemma 5.7. Then $c>0$. If there is a pair $(i_{2},j_{2})\in\mathcal{A}$ with $r_{i_{2}j_{2}}\geq 2$, then $\mathfrak{R}_{F}<1$ by Lemma 5.7 (1) and the inequality (7.1). So we may assume $r_{i_{2},j_{2}}=1$ for all $(i_{2},j_{2})\in\mathcal{A}$. If $c\geq 2$, then by Lemmas 4.14 and 5.7, we have $\mathfrak{R}_{F}<1$, which is impossible. Thus we have $c=1$ and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form $F_{1}^{\prime}$ of degree $d-1$. Since $l({\rm Aut}(F))$ contains $r_{i_{1}j_{1}}\geq 2$ and $1$, by Lemma 4.14 and Remark 4.15, we have $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ and $R(l(G),d)<11$, $2$, $11$, $6$ respectively. By Lemma 4.6, we have
@@
From this, we have $|H_{i_{1}j_{1}}|>60/11,360/2,25920/11,6531840/6$ if $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ respectively. On the other hand, $H_{i_{1}j_{1}}$ admitting an $F_{i_{1}j_{1}}$-lifting implies that $H_{i_{1}j_{1}}$ satisfies Lemma 6.2 (1)-(3), which is impossible by Table 1 (see e.g. [CCNPW] for character tables of simple groups). ∎
@@ 7.3. A stronger form of the main theorem
Let $X_{F}\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of degree $d\geq 3$ defined by $F$. If $n+2\geq 71$, then ${\rm J}(n+2)=(n+3)!$ by [Col07, Theorem A]. On the other hand, by Theorem 3.4, the optimal upper bound for the size of an abelian subgroup in ${\rm Aut}(F)$ is $d^{n+2}$. Thus the product ${\rm J}(n+2)\cdot d^{n+2}=d^{n+2}\cdot(n+3)!$ gives an upper bound for $|{\rm Aut}(F)|$, which gives an upper bound $d^{n+1}\cdot(n+3)!$ for $|{\rm Aut}(X_{F})|$. However, this bound for $|{\rm Aut}(X_{F})|$ is larger than the optimal bound $d^{n+1}\cdot(n+2)!=|{\rm Aut}(X_{d}^{n})|$.
@@
A smooth hypersurface $X$ in ${\mathbb{P}}^{n+1}$ of degree $d=1$ (resp. $2$) is isomorphic to ${\mathbb{P}}^{n}$ (resp. Fermat quadric $X_{2}^{n}$) with infinite automorphism group. Smooth cubic curves are elliptic curves and their automorphism groups are infinite. The automorphism group of a generic quartic surface is finite, but there are many smooth quartic surfaces with infinite automorphism groups. The automorphism group of Fermat quartic $X_{4}^{2}$ is infinite, and its finite subgroup ${\rm Lin}(X_{4}^{2})\cong C_{4}^{3}\rtimes S_{4}$ determines $X_{4}^{2}$ uniquely among all K3 surfaces ([Ogu05]).

### v2

@@ 7. Proof of the main theorem
In this section, using Theorem 6.1 and the refined bounds in Section 5, we classify smooth forms with large imprimitive or reducible automorphism groups in §7.1 and §7.2 (Theorems 7.1 and 7.4). In §7.3, combining these results, we prove a stronger form (Theorem 7.5) of Theorem 1.1.
@@ 7.1. Imprimitive ${\rm Aut}(F)$
Since ${\rm Aut}(F)$ is imprimitive and $F$ is not isomorphic to the Fermat form, $l({\rm Aut}(F))$ is of exponential type $r^{k}$ with $r\geq 2$ and $k\geq 2$, and ${\rm Aut}(F)$ has a primitive-constituent sequence $\mathcal{H}:=(H_{1},\ldots,H_{k})$ with $H_{i}\subset{\rm PGL}(r,{\mathbb{C}})$ and $H_{i}\cong H_{j}$ for $1\leq i,j\leq k$. By Lemma 4.7, we have $R(\mathcal{H},d)\geq 1$. By direct calculation (see (4.1)), we get $|H_{i}|>d^{r-1}\cdot r!$. In fact, $R(\mathcal{H},d)=\frac{d^{k}\cdot|H_{i}|^{k}\cdot k!}{d^{rk}\cdot(rk)!}\geq 1$ implies $|H_{i}|>d^{r-1}\cdot r!$ since $(rk)!>(r!)^{k}\cdot k!$. It suffices to consider two cases: $r\geq 3$ and $r=2$.
Case $r\geq 3$. If $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$, then by Theorem 6.1, we have $(r,d,|H_{1}|)=(3,4,168)$, $(3,6,360)$, $(4,4,1920)$, $(6,6,6531840)$. By direct calculation, we have $R(\mathcal{H},d)<1$, a contradiction. Thus, $H_{1}$ preserves no smooth forms of degree $d$. Since ${\rm JC}(r)>106$ for $r\geq 3$, by Lemmas 5.1, 5.5 and 4.15, we get $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$, a contradiction. In fact, we may write $F=F_{1}+F_{2}$, where $F_{1}\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ and $F_{2}\in(x_{r+1},\ldots,x_{n+2})\cdot{\mathbb{C}}[x_{1},\ldots,x_{n+2}]$ are forms of degree $d$ with $H_{1}$ preserving $F_{1}$. Then $F_{1}$ is not smooth and by Lemma 5.1, there exist integers $d_{1},\ldots,d_{r}\geq 0$ and a monomial $\mathfrak{m}\in F$ with $\mathfrak{m}=x_{1}^{d_{1}}\cdots x_{r}^{d_{r}}\cdot x_{r+j}$ for some $1\leq j\leq n+2-r$. Thus, by Lemmas 5.5 and 4.15 (1), we have
$|{\rm Aut}(F)|\leq\frac{R(l({\rm Aut}(F)),d)\cdot d^{n+2}\cdot(n+2)!}{{\rm JC}(r)}<\frac{106\cdot d^{n+2}\cdot(n+2)!}{{\rm JC}(r)}<d^{n+2}\cdot(n+2)!.$
Case $r=2$. If $H_{1}$ preserves no smooth forms of degree $d$, then $|{\rm Aut}(F)|<d^{n+2}\cdot(n+2)!$ by similar arguments as in case $r\geq 3$. Thus, we may assume $H_{1}$ preserves a smooth form $F_{1}(x_{1},\ldots,x_{r})$ of degree $d$. By Lemma 6.4, if $H_{1}\cong A_{4},S_{4},A_{5}$, then $H_{1}$ preserves no smooth forms in $2$ variables of degree less than $4,6,12$ respectively. If $H_{1}\cong A_{4}$, then $d\geq 4$ and $R(\mathcal{H},d)<1$, a contradiction. If $H_{1}\cong S_{4}$, then $d\geq 6$ and $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,6)$. Then $R(\mathcal{H},d)=4/3$ and $F$ can be expressed as $F_{1}(x_{1},x_{2})+F_{2}(x_{3},x_{4})+F^{\prime}$, where $F_{i}$ are preserved by $H_{i}$ and $F^{\prime}$ is in the intersection of the two ideals $(x_{1},x_{2})\cdot{\mathbb{C}}[x_{1},x_{2},x_{3},x_{4}]$ and $(x_{3},x_{4})\cdot{\mathbb{C}}[x_{1},x_{2},x_{3},x_{4}]$. If $F^{\prime}\neq 0$, then by Lemma 5.6, $H_{1}$ has a normal subgroup $\bar{H}_{1}$ and $|{\rm Aut}(F)|\leq\frac{B({\rm Aut}(F),F)}{[H_{1}:\bar{\bar{H}}_{1}]}$, where $\bar{\bar{H}}_{1}$ is a normal subgroup of $\bar{H}_{1}$ and $\bar{\bar{H}}_{1}$ preserves a non-zero form of degree $j$ for some $j\in\{1,2,3,4,5\}$. Then
$1\leq\frac{|{\rm Aut}(F)|}{d^{n+2}\cdot(n+2)!}\leq\frac{B({\rm Aut}(F),F)}{d^{n+2}\cdot(n+2)!\cdot[H_{1}:\bar{\bar{H}}_{1}]}=\frac{R(\mathcal{H},d)}{[H_{1}:\bar{\bar{H}}_{1}]}=\frac{4}{3\cdot[H_{1}:\bar{\bar{H}}_{1}]}$
and $\bar{\bar{H}}_{1}=H_{1}\cong S_{4}$, which contradicts Lemma 6.4. Thus, $F^{\prime}=0$. Therefore, $F$ is unique up to isomorphism by Lemma 6.4. If $H_{1}\cong A_{5}$, then $d=12$ or $d\geq 20$ by Lemma 6.4. So $R(\mathcal{H},d)\geq 1$ implies $(k,d)=(2,12),(3,12)$. Similar to the case $H_{1}=S_{4}$, we have the uniqueness of $F$. ∎
@@ 7.2. Reducible ${\rm Aut}(F)$
By the formula above, if $n_{m}=1$ or $m\geq 3$ or $n_{1}>2$, then all fractions on the right side are less than $1$, a contradiction to $q\geq 1$. Thus, we have $m=2$, $n_{1}=n_{2}=2$ and $q=\frac{25}{24}$. ∎
@@
Let $F=F(x_{1},\ldots,x_{r})$ be a smooth form of degree $d\geq 3$, where $r\geq 3$. If ${\rm Aut}(F)$ is reducible, then $|{\rm Aut}(F)|<d^{r}\cdot r!.$
@@
To prove the theorem by contradiction, we suppose $\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\geq 1$. Let $G:={\rm Aut}(F)$. By Theorem 3.9, we may assume at least one subdegree of $G$ is larger than $1$. Let $V_{i},W_{ij},k_{i},r_{ij},H_{ij},m,F_{ij}$ be as in $\bf{Set\textrm{-}up}$ 5.3. Recall that ${\mathbb{C}}^{r}=V_{1}\oplus\cdots\oplus V_{m}$, $V_{i}=W_{i1}\oplus\cdots\oplus W_{ik_{i}}$, $H_{ij}\subset{\rm PGL}(W_{ij})$, and ${\rm dim}(W_{ij})=r_{ij}$. We define $\beta$ to be $1$ (resp. $2$) if $l(G)$ has at least two different subdegrees (resp. all subdegrees of $G$ are the same). From the definitions, we have
@@
for any pair $(i,j)$ and any $1\leq k\leq k_{i}$. By Lemmas 4.7 and 4.15, we obtain $R(l(G),d)<60\beta$. If there is a monomial
@@
Case (1). Note that $V_{i}$ $(i=1,\ldots,m_{1})$ are irreducible $G$-stable subspaces of dimension $r_{i}:={\rm dim}(V_{i})=r_{i1}k_{i}$ and the irreducible linear groups $G_{i}:=p_{i}(G)$ preserve the smooth forms $F_{i}$ of degree $d$. Here $p_{i}:G\rightarrow{\rm GL}(V_{i})$ is given by $g\mapsto g|V_{i}$. Let $V^{\prime}:=\oplus_{i>m_{1}}V_{i}$ and $p^{\prime}:G\rightarrow{\rm GL}(V^{\prime})$ be given by $g\mapsto g|V^{\prime}$. Let $G_{m_{1}+1}:=p^{\prime}(G)$. Then $r=r_{1}+\cdots r_{m_{1}}+r_{m_{1}+1}$, where $r_{m_{1}+1}:={\rm dim}(V^{\prime})$. By Proposition 7.2, we have $\mathfrak{R}_{F_{i}}=\frac{|{\rm Aut}({F_{i}})|}{d^{r_{i}}\cdot r_{i}!}\leq 5/2$ $(i=1,2,\ldots,m_{1})$. Clearly, $G_{i}\subset{\rm Aut}({F_{i}})$ $(i=1,2,\ldots,m_{1}+1)$. Since $G_{m_{1}+1}$ consists of semi-permutations by $r_{ij}=1$ for $i>m_{1}$, by slightly adapting Theorem 3.9, we have $\mathcal{Q}:=\frac{|G_{m_{1}+1}|}{d^{r_{m_{1}+1}}\cdot r_{m_{1}+1}!}\leq 1$. Then by Lemma 7.3 and
$1\leq\mathfrak{R}_{F}=\frac{|{\rm Aut}(F)|}{d^{r}\cdot r!}\leq\frac{|G_{1}|\cdot|G_{2}|\cdots|G_{m_{1}}|\cdot|G_{m_{1}+1}|}{d^{r}\cdot r!}\leq\frac{(\prod_{i=1}^{m_{1}}\mathfrak{R}_{F_{i}}\cdot r_{i}!)\cdot(\mathcal{Q}\cdot r_{m_{1}+1}!)}{(r_{1}+\cdots+r_{m_{1}+1})!},$
@@
If $F^{\prime}$ is zero, then ${\rm Aut}(F)$ is irreducible (see Theorem 7.1), a contradiction. If $F^{\prime}\neq 0$, then by Lemma 5.6, $H_{11}$ has a normal subgroup $\bar{H}_{11}$ and $|{\rm Aut}(F)|\leq\frac{R(l(G),12)\cdot|\bar{\bar{H}}_{11}|\cdot d^{r}\cdot r!}{{\rm JC}(2)}=\frac{5\cdot|\bar{\bar{H}}_{11}|\cdot d^{r}\cdot r!}{144}$, where $\bar{\bar{H}}_{11}$ is a normal subgroup of $\bar{H}_{11}$ and $\bar{\bar{H}}_{11}$ preserves a non-zero form of degree $j$ for some $1\leq j\leq 11$. Since $H_{11}\cong A_{5}$ is a simple group, we have $\bar{\bar{H}}_{11}$ is either $H_{11}$ or trivial. From this and Lemma 6.4, we infer that $\bar{\bar{H}}_{11}$ is trivial and $\mathfrak{R}_{F}\leq\frac{5}{144}<1$, a contradiction. This proves case (1).
Case (2). Let $\mathcal{A}$ and $c$ be as in Lemma 5.7. Recall that
$\mathcal{A}=\{(i_{2},j_{2})|\;(i_{2},j_{2})\neq(i_{1},j_{1}),\;\exists\text{ amonomial }\mathfrak{m}\in S^{d-1}(W_{i_{1}j_{1}}^{*})\otimes S^{1}(W_{i_{2}j_{2}}^{*})\text{ in }F\}$
and $c$ is the cardinality of $\mathcal{A}$. Then $c>0$. If there is a pair $(i_{2},j_{2})\in\mathcal{A}$ with $r_{i_{2}j_{2}}\geq 2$, then $\mathfrak{R}_{F}<1$ by Lemma 5.7 (1) and the inequality (7.1). So we may assume $r_{i_{2},j_{2}}=1$ for all $(i_{2},j_{2})\in\mathcal{A}$. If $c\geq 2$, then by Lemmas 4.15 (3) and 5.7 (2), we have $R(l(G),d)<3$ and $\mathfrak{R}_{F}\leq\frac{R(l(G),d)}{d^{c-1}}<1$, which is impossible. Thus we have $c=1$ and $H_{i_{1}j_{1}}$ admits an $F_{i_{1}j_{1}}$-lifting preserving some non-zero form $F_{1}^{\prime}$ of degree $d-1$. Since $l({\rm Aut}(F))$ contains $r_{i_{1}j_{1}}\geq 2$ and $1$, by Lemma 4.15 (3) and Remark 4.16, we have $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ and $R(l(G),d)<11$, $2$, $11$, $6$ respectively. By Lemma 4.7, we have
@@
From this, we have $|H_{i_{1}j_{1}}|>60/11$, $360/2$, $25920/11$, $6531840/6$ if $r_{i_{1}j_{1}}=2$, $3$, $4$, $6$ respectively. On the other hand, $H_{i_{1}j_{1}}$ admitting an $F_{i_{1}j_{1}}$-lifting implies that $H_{i_{1}j_{1}}$ satisfies Lemma 6.2 (1)-(3), which is impossible by Table 1 (see [CCNPW] for character tables of finite simple groups). ∎
@@ 7.3. A stronger form of the main theorem
Let $X_{F}\subset{\mathbb{P}}^{n+1}$ be a smooth hypersurface of degree $d\geq 3$ defined by $F$. If $n+2\geq 71$, then ${\rm J}(n+2)=(n+3)!$ by [Col07, Theorem A] (note that this bound can be achieved by $S_{n+3}$). On the other hand, by Theorem 3.5, the optimal upper bound for the size of an abelian subgroup in ${\rm Aut}(F)$ is $d^{n+2}$. Thus the product ${\rm J}(n+2)\cdot d^{n+2}=d^{n+2}\cdot(n+3)!$ gives an upper bound for $|{\rm Aut}(F)|$, which gives an upper bound $d^{n+1}\cdot(n+3)!$ for $|{\rm Aut}(X_{F})|$. However, this bound for $|{\rm Aut}(X_{F})|$ is larger than the optimal bound $d^{n+1}\cdot(n+2)!=|{\rm Aut}(X_{d}^{n})|$.
@@
By Lefschetz hyperplane theorem, the Picard group of a smooth hypersurface $X$ of dimension at least $3$ is the same as for projective space. Hence ${\rm Aut}(X)={\rm Lin}(X)$. A smooth hypersurface $X$ in ${\mathbb{P}}^{n+1}$ of degree $d=1$ (resp. $2$) is isomorphic to ${\mathbb{P}}^{n}$ (resp. Fermat quadric $X_{2}^{n}$) with infinite automorphism group. Smooth cubic curves are elliptic curves and their automorphism groups are infinite. The automorphism group of a generic quartic surface is finite, but there are many smooth quartic surfaces with infinite automorphism groups. The automorphism group of Fermat quartic $X_{4}^{2}$ is infinite, and its finite subgroup ${\rm Lin}(X_{4}^{2})\cong C_{4}^{3}\rtimes S_{4}$ determines $X_{4}^{2}$ uniquely among all K3 surfaces ([Ogu05, Theorem 1.2 (i)]).